# Unconditional correctness of recent quantum algorithms for factoring and computing discrete logarithms

### v1

@@ 1.1. Context and quantum computing results
Public key cryptography has become a crucial element of our global digital communication infrastructure. Notable examples include the Diffie-Hellman key exchange [6] and the RSA (Rivest-Shamir-Adleman) cryptosystem [19], which rely on the difficulty of finding discrete logarithms and factoring large numbers, respectively.
However, in 1994, Peter Shor [20] developed an algorithm capable of efficiently solving these problems using a quantum computer.
Theorem (Shor [^1] The version of Shor’s algorithm stated here incorporates some small improvements from [21] (bounded number of calls) and [10] (fast integer multiplication). The number of qubits in Shor’s algorithm can be brought down to $O(n)$, though this typically requires a larger number of quantum gates (see e.g. [8, 12]).).
@@
[…]
@@
Shor’s original article [20] also includes a similar algorithm to solve the discrete logarithm problem. This advancement prompted the development of post-quantum cryptography, such as lattice-based cryptography, to ensure the security of our communications in the face of potential quantum computing breakthroughs [4].
Recently, Regev [18] devised a multidimensional variant of Shor’s factoring algorithm that reduces the size of the quantum circuit, i.e. the number of quantum gates, to $O(n^{3/2}\log n)$. A distinctive feature of Regev’s algorithm is that the quantum circuit must be called $O(\sqrt{n})$ times, rather than a constant number of times for Shor’s algorithm. This is not considered to be a serious drawback as the various complexity parameters of the quantum circuit are far more relevant metrics in quantum computing.
@@
First, the original version of Regev’s algorithm requires $O(n^{3/2})$ qubits – many more than Shor’s algorithm. The reason for this ultimately lies in the difficulty of performing classical computations on quantum computers in a space-efficient manner, due to the need for any quantum computation to be reversible. In a subsequent paper, Ragavan and Vaikuntanathan [17] improved Regev’s algorithm to use only $O(n\log n)$ qubits, while maintaining the circuit size at $O(n^{3/2}\log n)$ quantum gates. This work, inspired by ideas of Kalinski [13], essentially matches the space cost of Shor’s algorithm.
The second issue, which we address in this paper, is that Regev’s algorithm [18] does not have theoretical guarantees, unlike Shor’s algorithm. The correctness of Regev’s algorithm is based on an ad hoc number-theoretic conjecture which we describe below. This unproven assumption cannot be avoided as it lies at the core of Regev’s improvement on the circuit size. The variant of Ragavan and Vaikuntanathan [17] also crucially relies on this conjecture.
In this paper, we prove a version of Regev’s conjecture. This allows us to unconditionally prove the correctness of (slightly modified versions of) the algorithms of Regev [18] and Ragavan and Vaikuntanathan [17]. More precisely, we obtain the following algorithmic result.
@@
[…]
@@
This unconditionally establishes the results in [17, 18] up to logarithmic factors.
Regev’s algorithm was adapted by Ekerå and Gärtner [7] to the discrete logarithm problem. Their paper [7] also uses the space-saving arithmetic of Ragavan and Vaikuntanathan [17] to obtain a quantum circuit with $O(n^{3/2}\log n)$ gates and $O(n\log n)$ qubits for computing discrete logarithms.
Once more, the correctness of the algorithm by Ekerå and Gärtner [7] relies on an unproven hypothesis, which can be viewed as a stronger form of Regev’s conjecture. Our methods also apply to this stronger statement (again, with minor technical adjustments). Thus, we get an analogue of Theorem 1.1 for the discrete logarithm problem, i.e. an unconditional proof of the results in [7] up to logarithmic factors.
@@
[…]
@@
Moreover, the slight technical modifications that we need to introduce to make all these quantum algorithms unconditional are compatible with the error-correction results of [7, 17]. Finally, our results extend to further variants of Regev’s algorithm, such as computing discrete logarithms or multiplicative orders modulo $N$ for several elements simultaneously (see [7]).
@@
By being slightly more careful in the space usage of our quantum algorithms, it is possible to reduce the number of qubits to $O(n\log^{2}n)$ for Theorems 1.1 and 1.2. We will not show this here in order to reduce the quantum computing prerequisites to a minimum.
@@ 1.2. An overview of Regev’s algorithm
Let us start by recalling the basic idea behind Shor’s factoring algorithm [20].
Let $N$ be the integer to be factored, say odd and with at least two distinct prime factors. The first step is to generate a random integer $1<a<N$, which can be assumed to be coprime to $N$. Shor proved that there is an efficient quantum algorithm to find the multiplicative order $r$ of $a$ modulo $N$. It can be shown by elementary means that $a^{r/2}$ is a non-trivial square root of $1$ modulo $N$ (here we mean that $2\mid r$ and $a^{r/2}\not\equiv\pm 1$) with probability $\geqslant 1/2$. Whenever this is the case, we get a non-trivial divisor of $N$, namely $\gcd(N,a^{r/2}-1)$.
The reason that Shor’s algorithm uses $O(n^{2}\log n)$ quantum gates, where $n:=\lceil\log_{2}N\rceil$, comes from the part of the quantum circuit that performs modular exponentiation. Consider the task of computing a power $a^{M}\,(\mathrm{mod}\,N)$ on a classical computer, where $11<a<N-$ 1 and $M\leqslant N$. This computation can be performed efficiently using the well-known square-and-multiply method, which involves $O(\log M)$ multiplications of two integers modulo $N$. Since two $n$-bit integers can be multiplied in time $O(n\log n)$ by [10], the complexity of this modular exponentiation problem is thus $O(n^{2}\log n)$. [^2] Note that, even if $a$ is a small integer, say $a=2$, this procedure still takes time $O(n^{2}\log n)$, because most multiplications will involve $n$-bit integers. For Shor’s algorithm, a similar modular exponentiation needs to be performed quantumly, which requires $O(n^{2}\log n)$ quantum gates.
@@
The first idea in Regev’s algorithm is to work in higher dimensional space and replace the random parameter $a$ by several integers $b_{1},\ldots,b_{d}$ (chosen in a specific way, as we explain below). Eventually, the optimal choice of dimension turns out to be ${d\asymp\sqrt{\log N}}$. Similarly to Shor’s algorithm, the problem of factoring $N$ easily reduces to the task of finding a vector $(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}$ such that $\prod_{i=1}^{d}b_{i}^{e_{i}}$ is a non-trivial square root of $1$ modulo $N$.
Using additional tools such as the LLL lattice reduction algorithm, Regev [18] generalised Shor’s quantum algorithm to efficiently find all such vectors $(e_{1},\ldots,e_{d})$ in a ball of radius $N^{O(1/d)}$ centred at the origin.
@@
in the quantum setting, where the exponents $M_{1},\ldots,M_{d}$ are $\leqslant N^{O(1/d)}$. Regev observed that 1 can be computed classically in time $O(n^{3/2}\log n)$ if $|b_{i}|\leqslant(\log N)^{O(1)}$ for all $i$ (for $d\asymp\sqrt{n}\asymp\sqrt{\log N}$). This can be achieved by cleverly ordering the intermediate multiplications so that most of them involve integers with much fewer than $n$ bits. This classical procedure can then be turned into a quantum version that uses $O(n^{3/2}\log n)$ gates.
@@ 1.3. The number-theoretic conjecture behind Regev’s algorithm
Regev’s algorithm [18] and its space-efficient variant [17] crucially rely on the possibility of choosing very small integers $b_{1},\ldots,b_{d}$ such that $\prod_{i=1}^{d}b_{i}^{e_{i}}$ is a non-trivial square root of $1$ modulo $N$ for some $e_{1},\ldots,e_{d}$ with $|e_{i}|\leqslant N^{O(1/d)}$. We remind the reader that $n\asymp\log N$ and $d\asymp\sqrt{n}$.
The least restrictive bound [^3] The bound stated in Regev’s paper [18] is $|b_{i}|\leqslant(\log N)^{O(1)}$, but this condition can be relaxed somewhat, as observed by Ragavan (private communication). on the $b_{i}$’s that still allows for a quantum circuit with $\tilde{O}(n^{3/2})$ gates is to have $|b_{i}|\leqslant\exp(\tilde{O}(d))$, where the $\tilde{O}(\cdot)$ notation possibly hides a factor $(\log n)^{O(1)}$. Moreover, it is natural to set the $b_{i}$’s to be prime numbers (as in [7, 17, 18]) in order to avoid obvious multiplicative relations between them.
In this paper, we choose $b_{1},\ldots,b_{d}$ to be independent random prime numbers less than $d^{10^{3}d}$. With these parameters, we can now state a version of Regev’s conjecture [^4] Compare with [18, Theorem 1.1] or [17, Conjecture 3.1]. that follows from our results.
@@
The analogue of Regev’s algorithm for the discrete logarithm problem, proposed by Ekerå and Gärtner [7], relies on a stronger version of Regev’s conjecture. It concerns the geometry of a certain lattice $\mathcal{L}$ which encodes all multiplicative dependencies between the $\mathbf{b}_{i}$’s (modulo $N$). We prove the following version of it. [^5] Compare with [7, Assumption 1] or [17, Conjecture E.1]
@@
(2) :=$\mathcal{L}:=\Big{\{}(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}\ :\ \prod_{i=1}^{d}\mathbf{b}_{i}^{e_{i}}\equiv 1\!\!\pmod{N}\Big{\}}.$
@@
The main technical result of this paper is Theorem 2.18, of which Corollary 1.5 is a special case. Corollary 1.4 is a direct consequence Corollary 1.5.
@@ 1.4. Subgroup obstructions
Corollary 1.4 shows that, with high probability, every element $x$ in the subgroup generated by $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$ can be written as a short product of these $\mathbf{b}_{i}$ modulo $N$. To obtain the full strength of Regev’s assumption [17, Conjecture 3.1], it would be necessary to show that this subgroup $\mathopen{}\mathclose{{}\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}\right\rangle$ contains a non-trivial square root of $1$ modulo $N$.
Unfortunately, it is not possible to prove this in full generality without considerable advances on the well-known least quadratic non-residue problem in number theory. For any prime number $p$, write $n(p)$ for the smallest positive integer $a$ which is a quadratic non-residue (i.e. not a square) modulo $p$. The best known asymptotic upper bound for $n(p)$ is Burgess’s classical result [5] that
@@
    Assuming the Generalised Riemann Hypothesis, Ankeny proved the much stronger bound $n(p)\ll(\log p)^{2}$ [2]. This suggests that quadratic residues may no longer be an obstruction in this case, and indeed, under GRH, it is straightforward to show that $\mathopen{}\mathclose{{}\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}\right\rangle$ contains a non-trivial square root of $1$ modulo $N$ with high probability. Assuming GRH would also considerably simplify the proof of Theorem 2.18 and remove a logarithmic factor for the gate and qubit costs in Theorems 1.1 and 1.2. However, in this paper we seek fully unconditional results.
@@
    Using a grand zero-density estimate [11, Theorem 1 (1.8)] and the Landau-Page theorem [15, Corollary 11.10], it is possible to prove the above result unconditionally for almost all $N$, which would already have very interesting algorithmic consequences. More precisely, it can be shown that there exists a set of exceptions $E\subset\mathbb{N}$ satisfying
    $\mathopen{}\mathclose{{}\left|E\cap[1,x]}\right|\ll\exp\!\big{(}(\log x)^{1/2+o(1)}\big{)}$
    for $x\geqslant 1$, such that for every odd $N\not\in E$ with at least two prime factors, the subgroup $\mathopen{}\mathclose{{}\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}\right\rangle$ (with $\mathbf{b}_{i}=\mathbf{b}_{i}(N)$ as defined earlier) contains a non-trivial square root of $1$ modulo $N$ with high probability. From this and Corollary 1.4, it would follow that Regev’s algorithm finds a non-trivial divisor of $N$ with high probability for almost all $N$.
@@
    In this paper, we follow a different approach to obtain a completely unconditional result that applies to all $N$, by slightly modifying the algorithm itself. While the key to the efficiency of Regev’s algorithm is that the $\mathbf{b}_{i}$’s are small, one can tolerate a bounded number of large $\mathbf{b}_{i}$’s. [^6] In fact, this observation was already needed to adapt Regev’s algorithm to the discrete logarithm problem [7]. We use this extra flexibility to overcome the subgroup obstructions. The simplest way to proceed is to allow for one of the parameters, say $\mathbf{b}_{1}$, to be uniformly distributed in $(\mathbb{Z}/N\mathbb{Z})^{\times}$. As with Shor’s algorithm, this ensures that the subgroup $\mathopen{}\mathclose{{}\left\langle\mathbf{b}_{1}}\right\rangle$, and hence also $\mathopen{}\mathclose{{}\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}\right\rangle$, contains a non-trivial square root of $1$ with probability $\gg 1$.
@@
In his paper [18], Regev proposed to select the parameters $b_{1},\ldots,b_{d}$ deterministically, for example by setting $b_{i}$ to be the $i$-th smallest prime for $1\leqslant i\leqslant d$. While such a deterministic choice of parameters is likely to work for almost all $N$, this is probably very difficult to prove.
@@
In this paper, the parameters $\mathbf{b}_{i}$ are chosen to be random primes (not dividing $N$) under a certain threshold $X$. Another natural choice would have been to let $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$ be i.i.d. random variables uniformly distributed in the set of all integers $\leqslant X$ coprime to $N$. The proofs in this paper would carry over to this framework if an analogue of Proposition 2.9 for short character sums $\sum_{n\leqslant x}\chi(n)$ were available. Assuming the Generalised Riemann Hypothesis, the work of Granville and Soundararajan [9, Theorem 2] gives a bound of the required strength. However, this approach does not seem sufficient to obtain an unconditional result (the bounds in [9] become too weak when $L(s,\chi)$ has zeroes close to the line $\operatorname{Re}s=1$).
This paper is organised as follows. Our main technical result is Theorem 2.18, which shows that lattices $\mathcal{L}$ similar to that in Corollary 1.5 have a basis of short vectors with high probability. Using simple geometry of numbers (see Section 2.5), we reduce this problem to estimating the number of lattice points in balls of growing radii. Unfortunately, we are unable to obtain a suitable lattice point count for $\mathcal{L}$ directly. We resolve this by considering a different lattice $\mathcal{L}_{M}$ from the start of the argument (using the lemmas in Section 2.2). In Section 2.3, we expand the lattice point count for $\mathcal{L}_{M}$ in terms of Dirichlet characters modulo $N$. This produces a main term, which can be estimated precisely, and an error term. The heart of the proof lies in using a zero-density estimate for Dirichlet characters modulo $N$ to bound this error term unconditionally. Finally, we prove our quantum algorithmic applications (Theorems 1.1 and 1.2) in Section 3.
Acknowledgements
2. Proof of the existence of a short lattice basis

### v2

@@ 1.1. Context and quantum computing results
Public key cryptography has become a crucial element of our global digital communication infrastructure. Notable examples include the Diffie-Hellman key exchange [DH] and the RSA (Rivest-Shamir-Adleman) cryptosystem [RSA], which rely on the difficulty of finding discrete logarithms and factoring large numbers, respectively.
However, in 1994, Peter Shor [shor] developed an algorithm capable of efficiently solving these problems using a quantum computer.
Theorem (Shor [^1] The version of Shor’s algorithm stated here incorporates some small improvements from [shorbis] (bounded number of calls) and [annalsmult] (fast integer multiplication). The number of qubits in Shor’s algorithm can be brought down to $O(n)$, though this typically requires a larger number of quantum gates (see e.g. [gid, fastmult]).).
@@
$\mathbf{Input:}\text{ a composite integer $N\leqslant 2^{n}$}$
$\mathbf{Output:}\text{ a non-trivial divisor of $N$}$
@@
Shor’s original article [shor] also includes a similar algorithm to solve the discrete logarithm problem. This advancement prompted the development of post-quantum cryptography, such as lattice-based cryptography, to ensure the security of our communications in the face of potential quantum computing breakthroughs [postquantum].
Recently, Regev [regev] devised a multidimensional variant of Shor’s factoring algorithm that reduces the size of the quantum circuit, i.e. the number of quantum gates, to $O(n^{3/2}\log n)$. A distinctive feature of Regev’s algorithm is that the quantum circuit must be called $O(\sqrt{n})$ times, rather than a constant number of times for Shor’s algorithm. This is not considered to be a serious drawback as the various complexity parameters of the quantum circuit are far more relevant metrics in quantum computing.
@@
First, the original version of Regev’s algorithm requires $O(n^{3/2})$ qubits – many more than Shor’s algorithm. The reason for this ultimately lies in the difficulty of performing classical computations on quantum computers in a space-efficient manner, due to the need for any quantum computation to be reversible. In a subsequent paper, Ragavan and Vaikuntanathan [MIT] improved Regev’s algorithm to use only $O(n\log n)$ qubits, while maintaining the circuit size at $O(n^{3/2}\log n)$ quantum gates. This work, inspired by ideas of Kalinski [kaliski], essentially matches the space cost of Shor’s algorithm.
The second issue, which we address in this paper, is that Regev’s algorithm [regev] does not have theoretical guarantees, unlike Shor’s algorithm. The correctness of Regev’s algorithm is based on an ad hoc number-theoretic conjecture which we describe below. This unproven assumption cannot be avoided as it lies at the core of Regev’s improvement on the circuit size. The variant of Ragavan and Vaikuntanathan [MIT] also crucially relies on this conjecture.
In this paper, we prove a version of Regev’s conjecture. This allows us to unconditionally prove the correctness of (slightly modified versions of) the algorithms of Regev [regev] and Ragavan and Vaikuntanathan [MIT]. More precisely, we obtain the following algorithmic result.
@@
$\mathbf{Input:}\text{ a composite integer $N\leqslant 2^{n}$}$
$\mathbf{Output:}\text{ a non-trivial divisor of $N$}$
@@
This unconditionally establishes the results in [MIT, regev] up to logarithmic factors.
Regev’s algorithm was adapted by Ekerå and Gärtner [discretelogs] to the discrete logarithm problem. Their paper [discretelogs] also uses the space-saving arithmetic of Ragavan and Vaikuntanathan [MIT] to obtain a quantum circuit with $O(n^{3/2}\log n)$ gates and $O(n\log n)$ qubits for computing discrete logarithms.
Once more, the correctness of the algorithm by Ekerå and Gärtner [discretelogs] relies on an unproven hypothesis, which can be viewed as a stronger form of Regev’s conjecture. Our methods also apply to this stronger statement (again, with minor technical adjustments). Thus, we get an analogue of Theorem˜1.1 for the discrete logarithm problem, i.e. an unconditional proof of the results in [discretelogs] up to logarithmic factors.
@@
$\mathbf{Output:}\text{ an integer $x$ such that $g^{x}\equiv y\pmod{N}$}$
@@
Moreover, the slight technical modifications that we need to introduce to make all these quantum algorithms unconditional are compatible with the error-correction results of [discretelogs, MIT]. Finally, our results extend to further variants of Regev’s algorithm, such as computing discrete logarithms or multiplicative orders modulo $N$ for several elements simultaneously (see [discretelogs]).
@@
By being slightly more careful in the space usage of our quantum algorithms, it is possible to reduce the number of qubits to $O(n\log^{2}n)$ for Theorems˜1.1 and 1.2. We will not show this here in order to keep the quantum computing prerequisites to a minimum.
@@ 1.2. An overview of Regev’s algorithm
Let us start by recalling the basic idea behind Shor’s factoring algorithm [shor].
Let $N$ be the integer to be factored, which can be assumed to be odd and with at least two distinct prime factors. Shor proved that there is an efficient quantum algorithm to find the multiplicative order of any invertible element modulo $N$. With this in hand, factoring $N$ is straightforward. The first step is to pick an integer $1<a<N$ uniformly at random. If $\gcd(a,N)>1$, the factoring task is trivial. Otherwise, with probability at least $1/2$, the order of $a$ modulo $N$, denoted $r$, is even and satisfies $a^{r/2}\not\equiv\pm 1$ (i.e. $a^{r/2}$ is a non-trivial square root of $1$ modulo $N$). Whenever this is the case, we get a non-trivial divisor of $N$, namely $\gcd(N,a^{r/2}-1)$.
The reason that Shor’s algorithm uses $O(n^{2}\log n)$ quantum gates, where $n:=\lceil\log_{2}N\rceil$, comes from the part of the quantum circuit that performs modular exponentiation. Consider the task of computing a power $a^{M}\,(\mathrm{mod}\,N)$ on a classical computer, where 1<a<N11<a<N-1 and $M\leqslant N$. This computation can be performed efficiently using the well-known square-and-multiply method, which involves $O(\log M)$ multiplications of two integers modulo $N$. Since two $n$-bit integers can be multiplied in time $O(n\log n)$ by [annalsmult], the complexity of this modular exponentiation problem is thus $O(n^{2}\log n)$. [^2] Note that, even if $a$ is a small integer, say $a=2$, this procedure still takes time $O(n^{2}\log n)$, because most multiplications will involve $n$-bit integers. For Shor’s algorithm, a similar modular exponentiation needs to be performed quantumly, which requires $O(n^{2}\log n)$ quantum gates.
@@
The first idea in Regev’s algorithm is to work in higher-dimensional space and replace the random parameter $a$ by several integers $b_{1},\ldots,b_{d}$ (chosen in a specific way, as we explain below). Eventually, the optimal choice of dimension turns out to be $d\asymp\sqrt{\log N}$}. Similarly to Shor’s algorithm, the problem of factoring $N$ easily reduces to the task of finding a vector $(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}$ such that $\prod_{i=1}^{d}b_{i}^{e_{i}}$ is a non-trivial square root of $1$ modulo $N$.
Using additional tools such as the LLL lattice reduction algorithm, Regev [regev] generalised Shor’s quantum algorithm to efficiently find all such vectors $(e_{1},\ldots,e_{d})$ in a ball of radius $N^{O(1/d)}$ centred at the origin.
@@
in the quantum setting, where the exponents $M_{1},\ldots,M_{d}$ are $\leqslant N^{O(1/d)}$. Regev observed that ˜1 can be computed classically in time $O(n^{3/2}\log n)$ if $|b_{i}|\leqslant(\log N)^{O(1)}$ for all $i$ (for $d\asymp\sqrt{n}\asymp\sqrt{\log N}$). This can be achieved by cleverly ordering the intermediate multiplications so that most of them involve integers with much fewer than $n$ bits. This classical procedure can then be turned into a quantum version that uses $O(n^{3/2}\log n)$ gates.
@@ 1.3. The number-theoretic conjecture behind Regev’s algorithm
Regev’s algorithm [regev] and its space-efficient variant [MIT] crucially rely on the possibility of choosing very small integers $b_{1},\ldots,b_{d}$ such that $\prod_{i=1}^{d}b_{i}^{e_{i}}$ is a non-trivial square root of $1$ modulo $N$ for some $e_{1},\ldots,e_{d}$ with $|e_{i}|\leqslant N^{O(1/d)}$. We remind the reader that $n\asymp\log N$ and $d\asymp\sqrt{n}$.
[…]
In this paper, we choose $b_{1},\ldots,b_{d}$ to be independent random prime numbers less than $d^{10^{3}d}$. With these parameters, we can now state a version of Regev’s conjecture [^4] Compare with [regev, Theorem 1.1] or [MIT, Conjecture 1]. that follows from our results.
@@
The analogue of Regev’s algorithm for the discrete logarithm problem, proposed by Ekerå and Gärtner [discretelogs], relies on a stronger version of Regev’s conjecture. It concerns the geometry of a certain lattice $\mathcal{L}$ which encodes all multiplicative dependencies between the $\mathbf{b}_{i}$’s (modulo $N$). We prove the following version of it. [^5] Compare with [discretelogs, Assumption 1], or Conjecture E.1 in the full version of [MIT].
@@
(2) $\mathcal{L}:=\Big\{(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}\ :\ \prod_{i=1}^{d}\mathbf{b}_{i}^{e_{i}}\equiv 1\!\!\pmod{N}\Big\}$.
@@
The main technical result of this paper is Theorem˜3.18, of which Corollary˜1.5 is a special case. In turn, Corollary˜1.4 is a direct consequence of Corollary˜1.5.
@@ 1.4. Subgroup obstructions
Let $N$, $d$ and $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$ be as in Corollary˜1.4. This corollary shows that, with high probability, every element $x$ in the subgroup generated by $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$ can be written as a short product of these $\mathbf{b}_{i}$ modulo $N$. To obtain the full strength of Regev’s assumption [MIT, Conjecture 1], it would be necessary to show that this subgroup $\mathopen{}\mathclose{{\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}}\right\rangle$ contains a non-trivial square root of $1$ modulo $N$.
Unfortunately, it is not possible to prove this in full generality without considerable advances on the well-known least quadratic non-residue problem in number theory. For any prime number $p$, write $n(p)$ for the smallest positive integer $a$ which is a quadratic non-residue (i.e. not a square) modulo $p$. The best known asymptotic upper bound for $n(p)$ is Burgess’s classical result [burgess] that
@@
    Assuming the Generalised Riemann Hypothesis, Ankeny proved the much stronger bound $n(p)\ll(\log p)^{2}$ [ankeny]. This suggests that quadratic residues may no longer be an obstruction in this case, and indeed, under GRH, it is straightforward to show that $\mathopen{}\mathclose{{\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}}\right\rangle$ contains a non-trivial square root of $1$ modulo $N$ with high probability. Assuming GRH would also considerably simplify the proof of Theorem˜3.18 and remove a logarithmic factor for the gate and qubit costs in Theorems˜1.1 and 1.2. However, in this paper we seek fully unconditional results.
@@
    It is possible to prove the following partial result unconditionally: for almost all odd $N$ with at least two distinct prime factors, the subgroup $\mathopen{}\mathclose{{\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}}\right\rangle$ contains a non-trivial square root of $1$ modulo $N$ with high probability (with $d=d(N)$ and $\mathbf{b}_{i}=\mathbf{b}_{i}(N)$ as in Corollary˜1.4). By Corollary˜1.4, this implies that Regev’s original algorithm (with these parameters $\mathbf{b}_{i}$) finds a non-trivial divisor of $N$ with high probability, for almost all $N$. More precisely, the set $E\subset\mathbb{N}$ of exceptional values of $N$ for which the above does not hold can be shown to satisfy
    $\mathopen{}\mathclose{{\left|E\cap[1,x]}}\right|\leqslant\exp\!\big((\log x)^{1/2+o(1)}\big)$
    for $x\geqslant 1$. This partial result can be proved using character sums, the Landau-Page theorem [MV, Corollary 11.10] and a grand zero-density estimate [jutila, Theorem 1 (1.8)], by means of a case analysis depending on the prime factorisation of $N$.
@@
    In the present work, we follow a different approach to obtain a completely unconditional result that applies to all $N$, by slightly modifying the algorithm itself. While the key to the efficiency of Regev’s algorithm is that the $\mathbf{b}_{i}$’s are small, one can tolerate a bounded number of large $\mathbf{b}_{i}$’s. [^6] In fact, this observation was already needed to adapt Regev’s algorithm to the discrete logarithm problem [discretelogs]. We use this extra flexibility to overcome the subgroup obstructions. The simplest way to proceed is to allow for one of the parameters, say $\mathbf{b}_{1}$, to be uniformly distributed in $(\mathbb{Z}/N\mathbb{Z})^{\times}$. As with Shor’s algorithm, this ensures that the subgroup $\mathopen{}\mathclose{{\left\langle\mathbf{b}_{1}}}\right\rangle$, and hence also $\mathopen{}\mathclose{{\left\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{d}}}\right\rangle$, contains a non-trivial square root of $1$ with probability $\gg 1$.
@@
In his paper [regev], Regev proposed to select the parameters $b_{1},\ldots,b_{d}$ deterministically, for example by setting $b_{i}$ to be the $i$-th smallest prime for $1\leqslant i\leqslant d$. While such a deterministic choice of parameters is likely to work for almost all $N$, this seems to be very difficult to prove.
@@
In this paper, the parameters $\mathbf{b}_{i}$ are chosen to be random primes (not dividing $N$) under a certain threshold $X$. Another natural choice would have been to let $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$ be i.i.d. random variables uniformly distributed in the set of all integers $\leqslant X$ coprime to $N$. The proofs in this paper would carry over to this framework if an analogue of Proposition˜3.9 for short character sums $\sum_{n\leqslant x}\chi(n)$ were available. Assuming the Generalised Riemann Hypothesis, the work of Granville and Soundararajan [GS, Theorem 2] gives a bound of the required strength. However, this approach does not seem sufficient to obtain an unconditional result (the bounds in [GS] become too weak when $L(s,\chi)$ has zeroes close to the line $\operatorname{Re}s=1$).
1.5. Structure of the paper
We begin in Section˜2 with a proof of a simple special case of our main theorem. This is intended to illustrate our overall strategy and motivate the need for the more technical approach required in the general case. The purpose of this section is to build intuition; it is not formally required for the main argument.
Our main technical result is Theorem˜3.18, which shows that lattices $\mathcal{L}$ similar to that in Corollary˜1.5 have a basis of short vectors with high probability. Using simple geometry of numbers (see Section˜3.4), we reduce this problem to estimating the number of lattice points in balls of growing radii. Unfortunately, we are unable to obtain a suitable lattice point count for $\mathcal{L}$ directly. We resolve this by considering a different lattice $\mathcal{L}_{M}$ from the start of the argument (using the lemmas in Section˜3.1). In Section˜3.2, we expand the lattice point count for $\mathcal{L}_{M}$ in terms of Dirichlet characters modulo $N$. This produces a main term, which can be estimated precisely, and an error term. The heart of the proof lies in using a zero-density estimate for Dirichlet characters modulo $N$ to bound this error term unconditionally. Finally, we prove our quantum algorithmic applications (Theorems˜1.1 and 1.2) in Section˜4.